OpenStudy (anonymous):

How to covert the grams of a compound or element to moles? How to convert moles ompound or element to grams?

sam (.sam.):

To convert from grams to moles, simply divide the one's that you don't want , and multiply the one's that you want. For example : Let's say you have 50 grams of Sodium and you want to convert it to moles, you divide by its molar mass and then multiply the moles. \[50\mathbb{~grams ~of ~Na} \times \frac{1~mole}{23~grams}\] You can see that the grams cancel out perfectly and leave you with mole, which what we've trying to find out. \[50\mathbb{~\cancel{grams} ~of ~Na} \times \frac{1~mole}{23~\cancel{grams}}=2.17moles\] The \(23\) comes from the molar mass of Sodium (Na), It means that you have 23 grams in one mole of Sodium and that's Always fixed for sodium. Every element have different molar masses. You can look up different molar masses of each elements in the periodic table. -------------------------------------------------------------- To convert from moles to grams, Again same thing, divide the one's that you don't want , and multiply the one's that you want. For example : Let's say you have 2 moles of Potassium Chloride and you want to convert it to grams, now you have to multiply by its molar mass and then divide the moles. \[2 \mathbb{~moles~ of ~KCl} \times \frac{39+35.5~grams}{1~mole}\] \[2 \mathbb{~\cancel{moles}~ of ~KCl} \times \frac{39+35.5~grams}{1~\cancel{mole}}=149~grams ~ of ~ KCl\] You should know that why we're using 1 mole here because in one mole of KCl 'compound' , you have 39+35.5 grams of KCl . ================================================= For your information, just divide the one's that you don't want , and multiply the one's that you want. Try it out yourself :)